Entering Yellowstone National Park costs $20 for a car and $15 for a motorcycle. On a typical day, 5950 cars or motorcycles enter and pay a total of $107,875. How many motorcycles enter on a typical day?
Answer by rfer(16322)

let x = cars=3725
let y = mc=2225
x+y=5950= x=5950-y
20x+15y=107875
20(5950=-y)+15y=107875
119000-20y+15y=107875
-5y=-11125
y=2225